OpenStudy (anonymous):

Write the expression as either the sine, cosine, or tangent of a single angle. cosine of pi divided by three times cosine of pi divided by five plus sine of pi divided by three times sine of pi divided by five.

OpenStudy (anonymous):

cos (pi/3) cos(pi/5) + sin(pi/3)sin(pi/5)

OpenStudy (anonymous):

when I solve it I get 0.913545458 but now I do not know what angle it can be for sin , cos ,tan

OpenStudy (cwrw238):

you use the identity cos(A-B) = cosAcosB + sinAsinB

OpenStudy (anonymous):

how do I do that

OpenStudy (cwrw238):

let A = pi/3 and B = pi/5 - do you see how we can write it as the cos of one angle? ( work from right to left)

OpenStudy (anonymous):

so cos(pi/3-pi/5) = cospi/3cospi/5 + sinpi/3sinpi/5

OpenStudy (cwrw238):

right

OpenStudy (anonymous):

on the right I get 0.913545458

OpenStudy (anonymous):

and on the left the same

OpenStudy (cwrw238):

- yoou dont need to find the numerical values you on;y need to write it as a cosine cos (pi/3 - pi/5) = cos (2pi/15)

OpenStudy (anonymous):

wait so is that all I have to write ?

OpenStudy (anonymous):

or do I have to continue solving?

OpenStudy (cwrw238):

no thats all you have written it as the cosine of one angle

OpenStudy (anonymous):

oh ! okay okay , thank you so much for your help

OpenStudy (cwrw238):

yw

OpenStudy (cwrw238):

the fact that you get 0.9135454 on both sides does serve as a check though - that we have the right answer